OpenStudy (anonymous):

Find a polar form of the equation x^2+y^2-8x=0. a.) r= sin theta b.) r=8 c.) r=8 sin theta d.) r=8 cos theta

ganeshie8 (ganeshie8):

plugin : x = rcos(theta) y = rsin(theta) in the given equaiton

OpenStudy (anonymous):

so the x will be cosine and the y will be sin?

ganeshie8 (ganeshie8):

x will be r*cos(theta) y will be r*sin(theta)

ganeshie8 (ganeshie8):

x^2+y^2-8x=0 plugin : x = rcos(theta) y = rsin(theta) [r*cos(theta) ]^2 + [r*sin(theta)]^2 - 8*rcos(theta) = 0

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

okay, thanks for helping.

ganeshie8 (ganeshie8):

widout latex, its hard to type those sines/cosines :|

ganeshie8 (ganeshie8):

u sure u got how to simplify ? :)

OpenStudy (anonymous):

im actually looking it up on google how to simplify this.

ganeshie8 (ganeshie8):

[r*cos(theta) ]^2 + [r*sin(theta)]^2 - 8*rcos(theta) = 0 r^2 [cos^2(theta) + sin^2(theta) ] - 8*rcos(theta) = 0 r^2[1] - 8*rcos(theta) = 0 r^2 = 8*rcos(theta) r = 8cos(theta)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

yeah, it does. I've never done this before so i needed some step-by-step instruction.

ganeshie8 (ganeshie8):

good :) these are easy. u just need to remember below : x = rcos(theta) y = rsin(theta)

ganeshie8 (ganeshie8):

they give u : x^2+y^2 = r^2

ganeshie8 (ganeshie8):

to change to polar, u just substitute them for x and y, and simplify watever u can. thats all.

OpenStudy (anonymous):

okay, that explains a lot. thank you so much.

OpenStudy (usukidoll):

remember x =rcostheta y =rsintheta and r^2 = x^2 + y^2 you may need them again in calc iv

OpenStudy (anonymous):

this is calc? this is suppose to be trig....

OpenStudy (usukidoll):

O___________________________________o for me I had it in calc 3 and 4... wow my school must be that behind D:

OpenStudy (anonymous):

well this is an online class for me, because i failed the trig at school so thats probably why i dont understand anything. but thank you for giving the advice.

ganeshie8 (ganeshie8):

trig and polar wont lev u ever.. it pops up in many places... so u better master these now when u are at it :) good luck !